# MGMCpaper.bib

### v2

@@ 4.1 Random sample splitting
Note that the solution $u=A^{-1}f$ in (4.2) can be interpreted as the most likely state in the sampling problem, i.e. $u=\text{argmax}_{\theta\in\mathbb{R}^{n}}\{p(\theta)\}=\text{argmin}_{\theta\in\mathbb{R}^{n}}\{H(\theta)\}$. As already pointed out in Section 3.3, the solution of (4.2) can be obtained by splitting the matrix $A=:M-N$ and iterating
@@
Under certain conditions on the splitting defined by the matrix $M$ (see e.g. [saad2003iterative] for details) the iterates $u^{(k)}$ converge to the solution of (4.2). The corresponding iteration for sampling from the distribution with density given by (4.1) is
@@
and the same matrix $X$ as in (4.4). As shown in [Fox.C_Parker_2017_AcceleratedGibbsSampling], the following two statements are equivalent:
@@
The deterministic equivalent of the MGMG update in Alg. 1 is a standard multigrid cycle. Crucially, both algorithms can be expressed in the form in (4.3) and (4.5) respectively with the same splitting matrix $M$. Hence, as already pointed out in [goodman1989multigrid], the convergence analysis of the Multigrid Monte Carlo algorithm proceeds along the same lines as the well-established analysis of the corresponding multigrid method (see e.g. [Hackbusch.W_1985_MultiGridMethodsApplications, Hackbusch.W_2016_IterativeSolutionLarge]). The crucial observation is that the same iteration matrix $X$ arises in the deterministic multigrid error iteration in (4.4) and in the Multigrid Monte Carlo equivalent in (4.6); the explicit form of this matrix will be discussed in Section 4.3. Before doing this, it is instructive to first show that the MGMC update in Alg. 1 indeed leaves the target distribution invariant.
@@ 4.2 Invariance of the target distribution under MGMC updates
The following theorem formalises the invariance argument that is already presented in [goodman1989multigrid].
@@
Next, we consider the coarse level correction, i.e. lines 9–14 in Alg. 1. We show that given a sample $\theta_{\ell}$ on $\mathbb{R}^{n_{\ell}}$ from the target distribution, partially resampling $\psi_{\ell-1}$ on $\mathbb{R}^{n_{\ell-1}}$ from the correct distribution and adding this to $\theta_{\ell}$ as $\theta_{\ell}+I_{\ell-1}^{\ell}\psi_{\ell-1}$ (line 14 of Alg. 1) does not change the distribution. This result is essentially due to Liu and Sabatti [Liu.J.S_Sabatti_2000_generalised, Theorem 3], where the result was presented more abstractly. Since the results and proof in [Liu.J.S_Sabatti_2000_generalised] may seem somewhat opaque to a broader audience in numerical analysis, we recall the main points here and replicate the proof in Appendix A.1 for completeness.
@@
[…]
@@
Proposition 4.3 ([Liu.J.S_Sabatti_2000_generalised, Theorem 3]).
@@
[…]
@@
4:     $\chi_{\ell-1}^{(m+1)}:=\mathrm{MGMC}_{\ell-1}(A_{\ell-1},\widehat{f}_{\ell-1},\chi_{\ell-1}^{(m)})$triangleright Recursive call to $\mathrm{MGMC}_{\ell-1}$
@@ 4.3 Convergence of MGMC
$S_{\ell}^{\rho}$ $:=(M^{\rho}_{\ell})^{-1}N_{\ell}^{\rho},\quad\text{with $\rho\in\{\text{pre},\text{post}\}$}\quad\text{for $\ell\geq 1$},$ (4.16)
$S_{0}^{\mathrm{coarse}}$ $:=(M^{\mathrm{coarse}}_{0})^{-1}N_{0}^{\mathrm{coarse}},$
where $M^{\mathrm{coarse}}_{0}$, $M^{\mathrm{pre}}_{\ell}$ and $M^{\mathrm{post}}_{\ell}$ are the (invertible) splitting matrices which define the random coarse-, pre- and post-smoother respectively with $A_{\ell}=M^{\mathrm{pre}}_{\ell}-N^{\mathrm{pre}}_{\ell}=M^{\mathrm{post}}_{\ell}-N^{\mathrm{post}}_{\ell}$ and $A_{0}=M^{\mathrm{coarse}}_{0}-N^{\mathrm{coarse}}_{0}$.
@@
[…]
@@
where $A<B$ (resp. $A\leq B$) if and only if $B-A$ is positive definite (resp. positive semidefinite).
@@
is satisfied for all $\ell\geq 1$ with some constant $C_{A}>0$ independent of the level $\ell$.
@@
Proof. See [Hackbusch.W_2016_IterativeSolutionLarge, Theorem 11.59].
@@
Since $A_{\ell}$ is symmetric positive definite, the condition $0<A_{\ell}\leq M_{\ell}$ is always met for the symmetric Gauss–Seidel iteration; see [Hackbusch.W_1985_MultiGridMethodsApplications, Section 6.2.4.3]. We will see later that this condition also holds when MGMC is applied to linear Bayesian inverse problems described in Section 2.2, provided we use the low-rank smoothers introduced in Section 3.4.
To analyse the convergence of MGMC, we need to go beyond the usual (deterministic) multigrid theory, because it implies directly only the convergence of the mean; compare (4.17) and (4.26a) below. To discuss convergence of higher moments in the sampling context, we need extra work. Since we consider Gaussian random variables, it suffices to consider the convergence of the covariance. We start by writing down an explicit expression for the MGMC update in Alg. 1.
@@
where the random variables $W_{L}^{(m)}$ for each step $m$ are i.i.d. and have the same distribution as $W_{\ell}$ in (4.20) for $\ell=L$.
@@
[…]
@@
Various norm bounds for $X_{L}$ are available in the multigrid literature. In our Theorem 4.3 we quote a result for the symmetric case from [Hackbusch.W_2016_IterativeSolutionLarge, Theorem 11.59]. See for example [Hackbusch.W_1985_MultiGridMethodsApplications] for other varieties.
The following theorem formalises the equivalence of the convergence of multigrid and MGMC. It can be interpreted as a variant of [Fox.C_Parker_2017_AcceleratedGibbsSampling, Theorem 1] where the latter holds for general matrix splittings.
@@
Let the assumptions of Theorem 4.10 be satisfied. Then, the Kullback–Leibler divergence of the distribution of $(\theta_{L}^{(m)})_{m\in\mathbb{N}}$ and the target distribution $\mathcal{N}(A_{L}^{-1}f_{L},A_{L}^{-1})$ converges to 00. In particular, the sequence $(\theta_{L}^{(m)})_{m\in\mathbb{N}}$ converges in distribution to a Gaussian random variable with mean $A_{L}^{-1}f_{L}$ and covariance $A_{L}^{-1}$.
@@
The results above allow us to analyse properties of the Monte Carlo estimator based on the sample generated by MGMC: As for any Markov chain, the states $\theta_{L}^{(0)},\theta_{L}^{(1)},\theta_{L}^{(2)},\dots,\theta_{L}^{(m)}$ generated by the MGMC update in Alg. 1 are not independent. To quantify this dependence we measure the autocorrelations of the chain for a particular quantity of interest $\mathcal{F}$ which maps each state to a real number. We limit our analysis to quantities of interest that depend linearly on the sample state. To this end, we consider the linear functional $\mathcal{F}$ defined in (2.5); discretisation of $\mathcal{F}$ leads to the matrix representation $F_{L}$ given in (2.16). We analyse the integrated autocorrelation time (IACT) of the observed quantity $F_{L}^{\top}\theta_{L}^{(m)}$ which is defined as
@@
In the statistics literature the number of generated states divided by the IACT is also known as “effective sample size” (ESS) which can be interpreted as a measure for the number of statistically independent realisations of the quantity of interest in the chain.
The following theorem relates the IACT to the multigrid convergence rate $||X_{L}||_{A_{L}}$.
@@
Let $\tau_{\mathrm{int},F_{L}}$ be the IACT we would obtain if we drew the initial sample from the target multivariate normal distribution $\mathcal{N}(\mu_{L},A_{L}^{-1})$ with mean $\mu_{L}=A_{L}^{-1}f_{L}$ and covariance $A_{L}^{-1}$. Then (4.27) becomes
$\tau_{\mathrm{int},F_{L}}=1+2\sum_{s=1}^{\infty}\frac{F_{L}^{\top}X_{L}^{s}A_{L}^{-1}F_{L}}{F_{L}^{\top}A_{L}^{-1}F_{L}}.$ (4.28)
Assume further that the initial sample $\theta_{L}^{(0)}$ is drawn from a distribution with covariance $\mathrm{Cov}(\theta_{L}^{(0)})$ that is bounded in the sense that $||A_{L}^{1/2}\mathrm{Cov}(\theta_{L}^{(0)})A_{L}^{1/2}||_{2}<C_{0}$ for some constant $C_{0}$ that is independent of $L$. Then under the same assumption as in Theorem 4.10
$\tau_{\mathrm{int},F_{L}}\leq\frac{1+||X_{L}||_{A_{L}}}{1-||X_{L}||_{A_{L}}}\qquad\text{and}\qquad\left|\tau_{\mathrm{int},F_{L}}^{(m)}-\tau_{\mathrm{int},F_{L}}\right|\leq C\frac{||X_{L}||^{2m+1}_{A_{L}}}{(1-||X_{L}||_{A_{L}})(1-||X_{L}||_{A_{L}}^{2m})}$ (4.29)
@@
Note that $||X_{L}||_{A_{L}}$ does not have to be exceptionally small to obtain IACTs of $\sim 1-10$: For $||X_{L}||_{A_{L}}=0.5$ we get $\tau_{\mathrm{int},F_{L}}=2$ and for $||X_{L}||_{A_{L}}=0.8$ we have $\tau_{\mathrm{int},F_{L}}=9$. Furthermore, the condition $||A_{L}^{1/2}\mathrm{Cov}(\theta_{L}^{(0)})A_{L}^{1/2}||_{2}<C_{0}$ is trivially satisfied if the initial sample is drawn from an infinitely narrow delta-distribution, i.e. if we set $\theta_{L}^{(0)}=\theta_{L,0}$ for some fixed vector $\theta_{L,0}$.
@@
for the quantity of interest $I(F_{L}):=\mathbb{E}[F_{L}^{\top}\theta_{L}]$ with $\theta_{L}\sim\mathcal{N}(A_{L}^{-1}f_{L},A_{L}^{-1})$. The root-mean-squared-error for the sum in (4.30) can be bounded by the standard Monte Carlo rate $\mathcal{O}(M^{-1/2})$, where it turns out that the implied constant can be made independent of $L$. More precisely, we have the following results:
@@
Let Assumption 2.1 and assumptions of Theorems 4.10 and 4.12 hold. Assume further that the initial sample $\theta_{L}^{(0)}$ is drawn from a distribution with mean $\mathbb{E}[\theta_{L}^{(0)}]$ that is bounded in the sense that $\|\mathbb{E}[A_{L}^{1/2}\theta_{L}^{(0)}]\|_{2}<C_{0}$ for some constant $C_{0}$ that is independent of $L$. Suppose that $f_{L}$ is given by $(f_{L})_{j}=\langle f,\phi_{j}^{L}\rangle_{H}$ for ${j=1,\dots,n_{L}}$ for some $f\in H$. Then, the root-mean-squared-error of $\tilde{I}_{M}(F_{L})$ in (4.30) can be bounded as
@@
The original work in [goodman1989multigrid], which introduces the Multigrid Monte Carlo method, presents heuristic arguments for the invariance of the distribution under MGMC updates and this is formalised in our Theorem 4.1. The results in our Theorem 4.9, Lemma 4.8, Corollary 4.11 and Theorem 4.12 are also derived in a less rigorous form in [goodman1989multigrid].
To make the relationship to the work in [Fox.C_Parker_2017_AcceleratedGibbsSampling] explicit, we now show that the MGMC update can be written as a matrix splitting method of the form in (4.5) if the multigrid method is convergent. For this first recall that according to Lemma 4.6 the MGMC update can be written as $\theta_{L}^{(m+1)}=X_{L}\theta_{L}^{(m)}+Y_{L}f_{L}+W^{(m)}_{L}$ (c.f. (4.23)) with $Y_{L}=(I-X_{L})A_{L}^{-1}$ and $W_{L}^{(m)}\sim\mathcal{N}(0,A_{L}^{-1}-X_{L}A_{L}^{-1}X_{L}^{\top})$. Eliminating $X_{L}=I-Y_{L}A_{L}$, (4.23) can be written as
@@
Under the assumptions of Proposition 4.5 we have $||X_{L}||_{A_{L}}<1$ and therefore $Y_{L}$ is invertible as Lemma B.1 shows.
@@
which is exactly the form in (4.5). With this, we can interpret our Theorem 4.9 as a special case of [Fox.C_Parker_2017_AcceleratedGibbsSampling, Theorem 1] for the splitting matrix $M_{L}^{\mathrm{MGMC}}$ defined by Alg. 1. Furthermore, our Lemma 4.8 can be seen as a special case of [Fox.C_Parker_2017_AcceleratedGibbsSampling, Corollary 3].
4.4 Application to linear Bayesian inverse problems

### v3

@@ 4.1 Random sampling based on matrix splittings
Note that the solution $u=A^{-1}f$ in (4.2) can be interpreted as the most likely state in the sampling problem, i.e. $u=\text{argmax}_{\theta\in\mathbb{R}^{n}}\{p(\theta)\}=\text{argmin}_{\theta\in\mathbb{R}^{n}}\{H(\theta)\}$. As already pointed out in Section 3.3, the solution of (4.2) can be obtained by splitting the matrix A=:MNA=:M-N and iterating
@@
Under certain conditions on the splitting defined by the matrix $M$ (see e.g. [undefaaq] for details) the iterates $u^{(k)}$ converge to the solution of (4.2). The corresponding iteration for sampling from the distribution with density given by (4.1) is
@@
and the same matrix $X$ as in (4.4). As shown in [undefae], the following two statements are equivalent:
@@
The deterministic equivalent of the MGMG update in Alg. 1 is a standard multigrid cycle. Crucially, both algorithms can be expressed in the form in (4.3) and (4.5) respectively with the same splitting matrix $M$. Hence, as already pointed out in [undef], the convergence analysis of the Multigrid Monte Carlo algorithm proceeds along the same lines as the well-established analysis of the corresponding multigrid method (see e.g. [undefaar, undefah]). The crucial observation is that the same iteration matrix $X$ arises in the deterministic multigrid error iteration in (4.4) and in the Multigrid Monte Carlo equivalent in (4.6); the explicit form of this matrix will be discussed in Section 4.3. Before doing this, it is instructive to first show that the MGMC update in Alg. 1 indeed leaves the target distribution invariant.
@@ 4.2 Invariance of the target distribution under MGMC updates
The following theorem formalises the invariance argument that is already presented in [undef].
@@
Next, we consider the coarse level correction, i.e. lines 9–14 in Alg. 1. We show that given a sample $\theta_{\ell}$ on $\mathbb{R}^{n_{\ell}}$ from the target distribution, partially resampling $\psi_{\ell-1}$ on $\mathbb{R}^{n_{\ell-1}}$ from the correct distribution and adding this to $\theta_{\ell}$ as $\theta_{\ell}+I_{\ell-1}^{\ell}\psi_{\ell-1}$ (line 14 of Alg. 1) does not change the distribution. This result is essentially due to Liu and Sabatti [undefaas, Theorem 3], where the result was presented more abstractly. Since the results and proof in [undefaas] may seem somewhat opaque to a broader audience in numerical analysis, we recall the main points here and replicate the proof in Appendix A.1 for completeness.
@@
for some non-negative function $h_{\ell}$ on $\mathbb{R}^{n_{\ell}}$ and with normalising constant $Z_{\ell}:=\int_{\mathbb{R}^{n_{\ell}}}h_{\ell}(\theta_{\ell})d\theta_{\ell}$. Define the conditional coarse level density
@@
Proposition 4.3 ([undefaas, Theorem 3]).
@@
and $Z_{\ell-1}(\theta_{\ell})$ is a normalisation constant which ensures that $\int p_{\ell-1}(\psi_{\ell-1}|\theta_{\ell})\;d\psi_{\ell-1}=1$.
@@
4:  $\chi_{\ell-1}^{(m+1)}:=\mathrm{MGMC}_{\ell-1}(A_{\ell-1},\widehat{f}_{\ell-1},\chi_{\ell-1}^{(m)})$ $\triangleright$ Recursive call to $\mathrm{MGMC}_{\ell-1}$
@@ 4.3 Convergence of MGMC
$S_{\ell}^{\rho}$ $=(M^{\rho}_{\ell})^{-1}N_{\ell}^{\rho},\quad\text{with $\rho\in\{\text{pre},\text{post}\}$}\quad\text{for $\ell\geq 1$},$ (4.16)
$S_{0}^{\mathrm{coarse}}$ $=(M^{\mathrm{coarse}}_{0})^{-1}N_{0}^{\mathrm{coarse}},$
where $M^{\mathrm{coarse}}_{0}$, $M^{\mathrm{pre}}_{\ell}$ and $M^{\mathrm{post}}_{\ell}$ are the (invertible) splitting matrices which define the random coarse-, pre- and post-smoother, resp., with $A_{\ell}=M^{\mathrm{pre}}_{\ell}-N^{\mathrm{pre}}_{\ell}=M^{\mathrm{post}}_{\ell}-N^{\mathrm{post}}_{\ell}$ and $A_{0}=M^{\mathrm{coarse}}_{0}-N^{\mathrm{coarse}}_{0}$.
@@
Hence, if $\|X_{L}\|=:\sigma<1$ holds for any consistent matrix norm $\|\cdot\|$, then the error converges exponentially, i.e. $\|u_{L}^{(m)}-A_{L}^{-1}f_{L}\|<\sigma^{m}\|u_{L}^{(0)}-A_{L}^{-1}f_{L}\|$. Classical multigrid theory (see e.g. [undefaar]) states that the uniform bound on $\|X_{L}\|$ can be proven if the following two properties hold:
@@
where $A<B$ (resp. $A\leq B$) if and only if BAB-A is positive definite (resp. positive semidefinite).
@@
is satisfied for all $\ell\in\mathbb{N}$ with some constant $C_{A}>0$ independent of the level $\ell$.
@@
Proof. See [undefah, Theorem 11.59].
@@
Since $A_{\ell}$ is symmetric positive definite, the condition $0<A_{\ell}\leq M_{\ell}$ is always met for the symmetric Gauss–Seidel iteration; see [undefaar, Section 6.2.4.3]. We will see later that this condition also holds when MGMC is applied to linear Bayesian inverse problems described in Section 2.2, provided we use the low-rank smoothers introduced in Section 3.4.
The usual (deterministic) multigrid theory directly implies the convergence of the mean; compare (4.17) and (4.26a) below. To discuss convergence of higher moments in the sampling context, we need extra work. Since we consider Gaussian random variables, it suffices to consider the convergence of the covariance. We start by writing down an explicit expression for the MGMC update in Alg. 1.
@@
where the random variables $W_{L}^{(m)}$ for each step $m$ are i.i.d. and distributed as in (4.20) for $\ell=L$.
@@
where $X_{L}$ is defined in (4.15a) - (4.15b). Furthermore, if $X_{L}\colon(\mathbb{R}^{n_{L}},\|\cdot\|)\to(\mathbb{R}^{n_{L}},\|\cdot\|)$ defines a contraction with respect to $\|\cdot\|$, then $\mathbb{E}[\theta_{L}^{(m)}]$ converges to the solution $u_{L}$ of $A_{L}u_{L}=f_{L}$ and $\mathrm{Cov}(\theta_{L}^{(m)})$ converges to $A_{L}^{-1}$.
@@
Various norm bounds for $X_{L}$ are available in the multigrid literature. In our Theorem 4.3 we quote a result for the symmetric case from [undefah, Theorem 11.59]. See for example [undefaar] for other varieties.
The following theorem formalises the equivalence of the convergence of multigrid and MGMC. It can be interpreted as a variant of [undefae, Theorem 1] where the latter holds for general matrix splittings.
@@
Let the assumptions of Theorem 4.12 be satisfied. Then, the Kullback–Leibler divergence of the distribution of $(\theta_{L}^{(m)})_{m\in\mathbb{N}}$ and the target distribution $\mathcal{N}(A_{L}^{-1}f_{L},A_{L}^{-1})$ converges to 0. In particular, the sequence $(\theta_{L}^{(m)})_{m\in\mathbb{N}}$ converges in distribution to a Gaussian random variable with mean $A_{L}^{-1}f_{L}$ and covariance $A_{L}^{-1}$.
@@
The results above allow us to analyse properties of the Monte Carlo estimator based on the sample generated by MGMC: As for any Markov chain, the states $\theta_{L}^{(0)},\theta_{L}^{(1)},\theta_{L}^{(2)},\dots,\theta_{L}^{(m)}$ generated by the MGMC update in Alg. 1 are not independent. To quantify this dependence we measure the autocorrelations of the chain for a particular quantity of interest $\mathcal{F}$ which maps each state to a real number. We limit our analysis to quantities of interest that depend linearly on the sample state, which includes many practically relevant quantities of interest, such as fluxes or integrals over parts of the domain. To this end, we consider the linear functional $\mathcal{F}$ defined in (2.5); discretisation of $\mathcal{F}$ leads to the matrix representation $F_{L}$ given in (2.16).
We first analyse the integrated autocorrelation time (IACT) of the observed quantity $F_{L}^{\top}\theta_{L}^{(m)}$ which is defined as
@@
In the statistics literature the number of generated states divided by the IACT is also known as “effective sample size” (ESS) which can be interpreted as a measure for the number of statistically independent realisations of the quantity of interest in the chain. The IACT in (4.27) converges to
$\tau^{(\infty)}_{\mathrm{int},F_{L}}=1+2\sum_{s=1}^{\infty}\frac{F_{L}^{\top}X_{L}^{s}A_{L}^{-1}F_{L}}{F_{L}^{\top}A_{L}^{-1}F_{L}}.$ (4.28)
at a rate determined by $\|X_{L}\|_{A_{L}}$, as shown in the following theorem. Furthermore, if the initial sample is already drawn from the target normal distribution, the IACT simplifies directly to (4.28).
@@
Assume that the initial sample $\theta_{L}^{(0)}$ is drawn from a distribution with covariance $\mathrm{Cov}(\theta_{L}^{(0)})$ that is bounded in the sense that $A_{L}^{1/2}\mathrm{Cov}(\theta_{L}^{(0)})A_{L}^{1/2}||_{2}<C_{0}$ for some constant $C_{0}$ that is independent of $L$. Then under the same assumption as in Theorem 4.12
[…]
@@
Instead, if the initial sample $\theta_{L}^{(0)}$ is drawn from the target multivariate normal distribution, then the corresponding $\tau_{\mathrm{int},F_{L}}^{(m)}$, in (4.27), can be expressed as (4.28), regardless of the value of $m$.
@@
Note that $X_{L}||_{A_{L}}$ does not have to be exceptionally small to obtain IACTs of $\sim 1-10$: For $X_{L}||_{A_{L}}=0.5$ we get $\tau^{(\infty)}_{\mathrm{int},F_{L}}=2$ and for $X_{L}||_{A_{L}}=0.8$ we have $\tau^{(\infty)}_{\mathrm{int},F_{L}}=9$. Furthermore, the condition $A_{L}^{1/2}\mathrm{Cov}(\theta_{L}^{(0)})A_{L}^{1/2}||_{2}<C_{0}$ is trivially satisfied if the initial sample is drawn from an infinitely narrow delta-distribution, i.e. if we set $\theta_{L}^{(0)}=\theta_{L,0}$ for some fixed vector $\theta_{L,0}$.
@@
for the quantity of interest $I(F_{L}):=\mathbb{E}[F_{L}^{\top}\theta_{L}]$ with $\theta_{L}\sim\mathcal{N}(A_{L}^{-1}f_{L},A_{L}^{-1})$. The root-mean-squared-error for the sum in (4.30) can be bounded by the standard Monte Carlo rate $\mathcal{O}(M^{-1/2})$ and the implied constant can be made independent of $L$. More precisely, we have the following results:
@@
Let Assumption 1 and assumptions of Theorems 4.12 and 4.14 hold. Assume further that the initial sample $\theta_{L}^{(0)}$ is drawn from a distribution with mean $\mathbb{E}[\theta_{L}^{(0)}]$ that is bounded in the sense that $\|\mathbb{E}[A_{L}^{1/2}\theta_{L}^{(0)}]\|_{2}<C_{0}$ for some constant $C_{0}$ that is independent of $L$. Suppose that $f_{L}$ is given by $(f_{L})_{j}=\langle f,\phi_{j}^{L}\rangle_{H}$ for $j=1,\dots,n_{L}$} for some $f\in H$. Then, the root-mean-squared-error of $\tilde{I}_{M}(F_{L})$ in (4.30) can be bounded as
@@
The original work in [undef], which introduces the Multigrid Monte Carlo method, shows the invariance of the distribution under MGMC updates and this is formalised in our Theorem 4.1. The results in our Theorem 4.11, Lemma 4.10, Corollary 4.13 and Theorem 4.14 are also derived in [undef]. Since the original work in [undef] is aimed at the theoretical physics community, we rewrite the proofs here in a language more familiar to mathematicians and statisticians.
To make the relationship to the work in [undefae] explicit, we now show that the MGMC update can be written as a matrix splitting method of the form in (4.5) if the multigrid method is convergent. For this first recall that according to Lemma 4.8 the MGMC update can be written as $\theta_{L}^{(m+1)}=X_{L}\theta_{L}^{(m)}+Y_{L}f_{L}+W^{(m)}_{L}$ (c.f. (4.23)) with $Y_{L}=(I-X_{L})A_{L}^{-1}$ and $W_{L}^{(m)}\sim\mathcal{N}(0,A_{L}^{-1}-X_{L}A_{L}^{-1}X_{L}^{\top})$. Eliminating $X_{L}=I-Y_{L}A_{L}$, (4.23) can be written as
@@
Under the assumptions of Proposition 4.7 we have $X_{L}||_{A_{L}}<1$ and therefore $Y_{L}$ is invertible as Lemma B.1 shows.
@@
which is exactly the form in (4.5). With this, we can interpret our Theorem 4.11 as a special case of [undefae, Theorem 1] for the splitting matrix $M_{L}^{\mathrm{MGMC}}$ defined by Alg. 1. Furthermore, our Lemma 4.10 can be seen as a special case of [undefae, Corollary 3].
4.4 Application to linear Gaussian Bayesian inverse problems